64^x=1/2 and x^4/3=625
Answer by AnlytcPhil(1116)

  64x = 1/2

  64x = 1/21

(26)x = 2-1

  26x = 2-1

   6x = -1

    x = -1/6

-----------------------

     x4/3 = 625

     x4/3 = 54

(x4/3)3/4 = (54)3/4

       x1 = 53

        x = 125

Edwin
AnlytcPhil@aol.com